OpenStudy (foreverandeverandeverrrr021014):

Let f(x) = x − 2 and g(x) = x2 − 7x − 9. Find f(g(−1)).

OpenStudy (boldjon):

coincidentally, g(-1) = -1, but the input and output are usually not the same

OpenStudy (boldjon):

so f(g(-1)) = -3

OpenStudy (boldjon):

well you start with f(x) = x-2 then you replace x with g(-1) since you want f(g(-1))

OpenStudy (boldjon):

so that explains how we got f(g(-1)) = g(-1) - 2 then on the right side you replace g(-1) with -1 because above we've shown that g(-1) = -1

OpenStudy (boldjon):

g(x) = x^2 - 6x - 7 g(2) = (2)^2 - 6(2) - 7 g(2) = ???

OpenStudy (boldjon):

g(2) = -15

OpenStudy (boldjon):

and that's yo answer

OpenStudy (dinamix):

f0g = f(g(x))= x^2-7x-11 = f(g(-1)) = 1+7-11= -3

OpenStudy (foreverandeverandeverrrr021014):

thank you soo much

OpenStudy (dinamix):

i 'm true 200% @boldjon